# Comparison of Generative Learning Methods for Turbulence Modeling

### v1

@@ 3.1 Variational Autoencoders (VAE)
where $p_{\theta}$ and $q_{\phi}$ are the probabilistic encoder and decoder respectively and $\mathfrak{d}_{KL}$ is the Kullback-Leibler divergence [77], which quantifies how different $q_{\phi}(z|x)$ is from $p_{\theta}(z)$. Note that the right-hand side of (1) shows the reconstruction loss and the left-hand term forces the latent distribution to be close to the Gaussian prior. Since we adopt the ubiquitous assumption of a standard normal prior $p_{\theta}(z)$ and Gaussian posterior $q_{\phi}(z|x)$, the KL-divergence can be calculated analytically and only the reconstruction loss needs to be estimated by sampling. Thus, we can rewrite the optimization problem in terms of
With our prior and posterior assumptions, we can simply sample a point $z$ from the standard normal distribution and run it through the trained probabilistic decoder at inference time. However, for sampling at training time, we need to use the so-called reparameterization trick (see figure 1). We want to optimize the distribution parameters $\mu$ and $\sigma$ of the latent distribution $\zeta$ from which we want to sample the inputs for the decoder. The problem is that we can’t backpropagate [78] through a stochastic sample because we can’t compute gradients for it. Hence, the idea of reparameterization is to treat $\mu$ and $\sigma$ as deterministic variables, sample a point $\varepsilon$ of the standard normal distribution, and define a new latent sample $z=\mu+\sigma\cdot\varepsilon$. With the help of this trick, the stochastic nature of the distribution parameters is disentangled so that we can compute the gradients of $\mu$ and $\sigma$ while still preserving the stochasticity, since $\bm{z}$ is also a random variable due to the randomness of $\varepsilon$. Note that since the probabilistic decoder’s inputs are stochastic, the original images are not only reconstructed as in classical autoencoders, but new images can be generated.
3.2 Deep Convolutional Generative Adversarial Networks (DCGAN)
Figure 2: Architecture of a deep convolutional GAN [24, 27]. While the training data is given by the real-world data $\mathcal{X}$, the fake samples $G(\bm{z})\sim G_{*}\lambda$ are produced by the generator from a noise vector $\bm{z}$. The disriminator’s inputs are GAN-synthesized and real-world samples, and it’s task is to estimate the probability in a range of $[0,1]$ that an input sample comes from $\mathcal{X}$ rather than being generated by $G$. During training, the feedback from the discriminator reaches the generator by backpropagation. The entire GAN framework can be backpropagated at once, since $G$ and $D$ are both fully differentiable and trained end-to-end. If the unknown distribution of the real-world data is approximated by $G$ (i.e. $G_{*}\lambda\approx\mu$) and $D$ is only able to guess the real-world from the fake samples (i.e. $D(\cdot)\approx\nicefrac{{1}}{{2}}$), the problem (3) reached its optimum.
Similar to AE, generative adversarial networks (GAN) are basically composed of two mappings - a generator $G:\Lambda\rightarrow\Omega$ and a discriminator $D:\Omega\rightarrow[0,1]$, where $\Lambda$ is the space of latent variables with an easily simulated probability measure $\lambda$ as Gaussian noise, $\Omega=\{\mathcal{X},\{G(z)\}\}$ is the space of real-world and generated images, and the interval $[0,1]$ gives the probability whether a sample is from the real-world images or a generated one. The generator $G$ transforms the noise measure $\lambda$ to the image measure $G_{*}\lambda$. The objective of adversarial learning is to train a mapping $G$ using feedback from the discriminator $D$, such that $D$ cannot differentiate between synthetic samples generated by $G_{*}\lambda$ and real samples drawn from the unknown target distribution $\mu$. The discriminator $D$, in turn, is trained as a classifier to assign high probabilities to real-world data and low probabilities to synthesized data. If the generator $G$ has been trained so effectively that even the most optimal discriminator $D$ is unable to distinguish between samples from $\mu$ and $G_{*}\lambda$, then the generative learning process is considered successful, see figure 2.
In practice, both the generator $G$ and the discriminator $D$ are neural networks. The feedback from $D$ to $G$ is propagated backward [78] through the composite mapping $D\circ G$, allowing the neural network weights of $G$ to be updated. Moreover, the universal approximation theorem for (deep) neural networks ensures that any mappings $G$ and $D$ can be approximated to a desired level of precision if the neural networks have a sufficiently wide and deep architecture. For qualitative and quantitative results, see [24, 79, 80, 81, 57, 82]. This work investigates the advanced deep convolutional GAN (DCGAN) [27] framework. As the name suggests, the generator $G$ and the discriminator $D$ are deep convolutional neural networks (DCNN), which have been successfully applied to image processing [83, 84]. For guidelines on how to properly integrate DCNN into GAN to ensure stable training at high resolution with deeper architecture, see [27].
The training of a GAN is structured as a minimax game between the discriminator $D$ and the generator $G$, which is mathematically represented by the min-max optimization problem
Note that the loss function $\mathcal{L}(D,G)$ is commonly known as binary cross-entropy [85]. As observed in [24], taking the maximum over a sufficiently large hypothesis space $\mathcal{H}_{D}$ of discriminators yields to
where, $\mathfrak{d}_{\text{JS}}(\mu\|G_{*}\lambda)$ is the Jensen-Shannon divergence [86], which is an information-theoretic pseudo-distance between the invariant measure $\mu$ and the generated measure $G_{*}\lambda$ and is defined as
$\mathfrak{d}_{\text{JS}}(\mu\|G_{*}\lambda)=\mathfrak{d}_{\text{KL}}\left(\mu\left\|\frac{G_{*}\lambda+\mu}{2}\right.\right)+\mathfrak{d}_{\text{KL}}\left(G_{*}\lambda\left\|\frac{G_{*}\lambda+\mu}{2}\right.\right)~{}.$ (5)
@@ 3.2 Deep Convolutional Generative Adversarial Networks (DCGAN)
Here, the Kulback-Leibler distance between the measures $\nu$ and $\mu$ is given by $\mathfrak{d}_{\text{KL}}(\mu\|\nu)=-\mathbb{E}_{\bm{x}\sim\mu}\left[\log\left(\frac{f_{\nu}}{f_{\mu}}(\bm{x})\right)\right]$ with continuous probability densities $f_{\mu}$ and $f_{\nu}$, respectively. Note that $\mathfrak{d}_{\text{KL}}(\mu\|\nu)=0$ is only true if and only if $f_{\mu}(\bm{x})=f_{\nu}(\bm{x})$ with $\mu$-probability one and hence $\mu=\nu$. As a consequence, $\mathfrak{d}_{\text{JS}}(\mu\|G_{*}\lambda)$ also measures the distance between $\mu$ and $G_{*}\lambda$.
@@
Note also that GAN have been shown to be a sound mathematical approach to turbulence modeling, as one can theoretically prove that they converge for ergodic systems. For more details, see our previous work [28].
@@ 3.3 Denoising Diffusion Probabilistic Models (DDPM)
Diffusion models convolve the distribution we are interested in into noise via an iterative Markov chain process [26, 25], by adding noise at each timestep. The model then learns the backwards map, that is, the function that returns the noise to the original distribution. The forwards process contains no learnable parameters, and the backwards process is modelled with a neural network, parameterised by weights $\theta$. Figure 3 gives an outline of the model architecture.
Figure 3: Denoising diffusion probabilistic (DDPM) model architecture [25]. At training time, a sample from the dataset $x\sim\Omega$ is noised to a random step $t$ in the noising process, producing a partially noised sample $x_{t}\sim\mathcal{N}(x_{t-1}\sqrt{1-\beta_{t}},\beta_{t}I)$. The UNet is trained to produce $x_{t-1}$ from $x_{t}$ with the parameter $t$ given as a positional embedding. At inference time, a normal sample $x_{T}=z\sim\mathcal{N}(0,1)$ is produced and fed to the UNet, which reverses the noising steps one at a time and produces a new sample $x^{\prime}\sim\Omega^{\prime}$.
@@ 3.3.1 VAE Interpretation
Broadly speaking, there are two schools of thought by which to characterise the theory of diffusion models. We will begin with the VAE interpretation, which DDPM also subscribes to. Adapting the minimisation problem 1 from the VAE section above gives
@@
Here $x_{t}$ represents the state of the datapoint $x_{0}$ at timestep $t$ in the Markov chain, which has $T$ total steps. $x_{T}$ is equivalent to the latent variable $z$, the fully noised datapoint, assumed to be drawn from a standard unit normal with dimension equal to that of the datapoint. The particular implementation used in this paper, denoising diffusion probabilistic models (DDPM) [25], parameterises the forward noising process with a normal distribution
$q(x_{t}|x_{t-1})=\mathcal{N}(x_{t};\,x_{t-1}\sqrt{1-\beta_{t}},\beta_{t}I)$ (7)
The $\{\beta_{t}\}$ represent a fixed noise schedule, often linear in $t$, chosen so that $q(x_{T}|x_{0})\approx\mathcal{N}(0,I)$ [87]. This gives rise to the backwards model, notably also a normal distribution, with analytical form
@@
We can do one better and give the forms of $\tilde{\mu}_{t}(x_{t},x_{0})$ and $\tilde{\beta}_{t}$. The full derivation can be found in [88],
@@
and make the choice that $\Sigma_{\theta}(x_{t},t)=\sigma_{t}^{2}I$. We are finally ready to derive the ELBO, as we did with the VAE. Inserting into the minimisation problem 6, and noting that the form of the term in inside the sum can be written [88, 26, 25]
@@
for some $C$ independent of the parameters $\theta$. By itself, this is suitable for optimisation, but we can go one step further, using the reparameterisation trick mentioned in 3.1 [23]. We separate the noise from the data to give
@@
for our minimisation objective. In practice, we also drop the prefactor since it only scales the loss. In [25] it is noted that, aside from being simpler, this empirically leads to better results.

### v2

@@ 3.1 Variational Autoencoders (VAE)
Here, $p_{\theta}(x|z)$ is the likelihood of $x$ given $z$, $p_{\theta}(z)$ is the prior, and $p_{\theta}(x)=\int p_{\theta}(x|z)p_{\theta}(z)dz$ is the evidence, the total probability of observing $x$ under the model, averaged over all possible latent configurations $z$, which is generally expensive and intractable to compute. The solution is to approximate the evidence with a family of distributions $q_{\phi}(z|x)$, where $\phi$ are the network weights for the distribution parameters for each data point $x$. Therefore, a computationally tractable approximation of the posterior can be derived by optimizing the evidence lower bound (ELBO), which leads to the following optimization problem for training the VAE:
where $p_{\theta}$ and $q_{\phi}$ are the probabilistic encoder and decoder respectively and $\mathfrak{d}_{KL}$ is the Kullback-Leibler (KL) divergence [kullback1951information], which quantifies the difference between two probability distributions by computing the expected logarithmic difference between them, i.e. how much information is lost when using $q_{\phi}(z|x)$ to approximate $p_{\theta}(z)$. Although it behaves like a distance between two distributions, it’s not a true norm, being asymmetric and not satisfying the triangle inequality. Note that the right-hand side of (1) shows the reconstruction loss and the left-hand term forces the latent distribution to be close to the Gaussian prior. Since we adopt the ubiquitous assumption of a standard normal prior $p_{\theta}(z)$ and Gaussian posterior $q_{\phi}(z|x)$, the KL-divergence can be calculated analytically and only the reconstruction loss needs to be estimated by sampling. Thus, we can rewrite the optimization problem in terms of
With our prior and posterior assumptions, we can simply sample a point $z$ from the standard normal distribution and run it through the trained probabilistic decoder at inference time. However, for sampling at training time, we need to use the so-called reparameterization trick (see figure˜1). We want to optimize the distribution parameters $\mu$ and $\sigma$ of the latent distribution $\zeta$ from which we want to sample the inputs for the decoder. The problem is that we can’t backpropagate [backprop] through a stochastic sample, i.e., we cannot compute gradients of the loss with respect to network parameters if $z$ is sampled directly. Hence, the idea of reparameterization is to treat $\mu$ and $\sigma$ as deterministic variables, sample a point $\varepsilon$ from the standard normal distribution, and define a new latent sample $z=\mu+\sigma\cdot\varepsilon$. This disentangles the stochasticity of the latent variable from the parameters, allowing gradients of $\mu$ and $\sigma$ to be computed while preserving randomness through $\varepsilon$. Importantly, since the decoder’s inputs are stochastic, it does not merely reconstruct the original images pixel-by-pixel, but can generate new samples within the same domain (e.g., turbulent flow data), capturing the variability learned during training.
3.2 Deep Convolutional Generative Adversarial Networks (DCGAN)
Figure 2: Architecture of a deep convolutional GAN [vanilla_gan, dcgan]. While the training data is given by the real-world data $\mathcal{X}$, the fake samples $G(\bm{z})\sim G_{*}\lambda$ are produced by the generator from a noise vector $\bm{z}$. The discriminator’s inputs are GAN-synthesized and real-world samples, and it’s task is to estimate the probability in a range of $[0,1]$ that an input sample comes from $\mathcal{X}$ rather than being generated by $G$. During training, the feedback from the discriminator reaches the generator by backpropagation. The entire GAN framework can be backpropagated at once, since $G$ and $D$ are both fully differentiable and trained end-to-end. If the unknown distribution of the real-world data is approximated by $G$ (i.e. $G_{*}\lambda\approx\mu$) and $D$ is only able to guess the real-world from the fake samples (i.e. $D(\cdot)\approx\frac{1}{2}$), the problem (3) reached its optimum.
Similar to AE, generative adversarial networks (GAN) are basically composed of two mappings - a generator $G:\Lambda\rightarrow\Omega$ and a discriminator $D:\Omega\rightarrow[0,1]$, where $\Lambda$ is the space of latent variables with an easily simulated probability measure $\lambda$ as Gaussian noise, $\Omega=\{\mathcal{X},\{G(z)\}\}$ is the space of real-world and generated images, and the interval $[0,1]$ gives the probability whether a sample is from the real-world images or a generated one. The generator $G$ transforms the noise measure $\lambda$ to the image measure $G_{*}\lambda$. The objective of adversarial learning is to train a mapping $G$ using feedback from the discriminator $D$, such that $D$ cannot differentiate between synthetic samples generated by $G_{*}\lambda$ and real samples drawn from the unknown target distribution $\mu$. The discriminator $D$, in turn, is trained as a classifier to assign high probabilities to real-world data and low probabilities to synthesized data. If the generator $G$ has been trained so effectively that even the most optimal discriminator $D$ is unable to distinguish between samples from $\mu$ and $G_{*}\lambda$, then the generative learning process is considered successful, see figure˜2.
In practice, both the generator $G$ and the discriminator $D$ are neural networks. The feedback from $D$ to $G$ is propagated backward [backprop] through the composite mapping $D\circ G$, allowing the neural network weights of $G$ to be updated. Moreover, the universal approximation theorem for (deep) neural networks ensures that any mappings $G$ and $D$ can be approximated to a desired level of precision if the neural networks have a sufficiently wide and deep architecture. For qualitative and quantitative results, see [vanilla_gan, asatryan2020convenient, pix2pixHD, CycleGAN2017, wang2018esrgan, karras2019style]. This work investigates the advanced deep convolutional GAN (DCGAN) [dcgan] framework. As the name suggests, the generator $G$ and the discriminator $D$ are deep convolutional neural networks (DCNN), which have been successfully applied to image processing [cnn_impact, cnn_explanation]. For guidelines on how to properly integrate DCNN into GAN to ensure stable training at high resolution with deeper architecture, see [dcgan].
The training of a GAN is structured as a minmax game between the discriminator $D$ and the generator $G$, which is mathematically represented by the min-max optimization problem
Note that the loss function $\mathcal{L}(D,G)$ is commonly known as binary cross-entropy [bce]. As observed in [vanilla_gan], taking the maximum over a sufficiently large hypothesis space $\mathcal{H}_{D}$ of discriminators yields to
where, $\mathfrak{d}_{\text{JS}}(\mu\|G_{*}\lambda)$ is the Jensen-Shannon (JS) divergence [menendez1997jensen], which is an information-theoretic pseudo-distance between the invariant measure $\mu$ and the generated measure $G_{*}\lambda$ and is defined as
@@ 3.2 Deep Convolutional Generative Adversarial Networks (DCGAN)
Here, the Kulback-Leibler distance between the measures $\nu$ and $\mu$ is given by $\mathfrak{d}_{\text{KL}}(\mu\|\nu)=-\mathbb{E}_{\bm{x}\sim\mu}\left[\log\left(\frac{f_{\nu}}{f_{\mu}}(\bm{x})\right)\right]$ with continuous probability densities $f_{\mu}$ and $f_{\nu}$, respectively. Note that $\mathfrak{d}_{\text{KL}}(\mu\|\nu)=0$ is only true if and only if $f_{\mu}(\bm{x})=f_{\nu}(\bm{x})$ with $\mu$-probability one and hence $\mu=\nu$. Consequently, $\mathfrak{d}_{\text{JS}}(\mu\|G_{*}\lambda)$ also measures the distance between $\mu$ and $G_{*}\lambda$, and, being symmetric and bounded, it provides a more stable measure of distributional difference than the KL divergence alone - a property exploited in GAN training.
@@
Note also that GAN have been shown to be a sound mathematical approach to turbulence modeling, as one can theoretically prove that they converge for ergodic systems[neumann1932proof, birkhoff1931proof], which are systems in which a single deterministic trajectory explores the entire phase space and thus allows for a statistical description of the system’s behavior [Buzzi2009, eisner2015operator]. For more details, see our previous work [drygala2022generative].
@@ 3.3 Denoising Diffusion Probabilistic Models (DDPM)
Diffusion models convolve the distribution we are interested in into noise via an iterative Markov chain process [sohldickstein2015deep, ho2020denoising], by adding noise at each timestep. The model then learns the backwards map, that is, the function that returns the noise to the original distribution. In other words, the forwards process progressively corrupts the data with noise, while the backwards process denoises it to reconstruct the original distribution. The forwards process contains no learnable parameters, and the backwards process is modeled with a neural network, parameterized by weights $\theta$. Figure 3 gives an outline of the model architecture.
Figure 3: Denoising diffusion probabilistic (DDPM) model architecture [ho2020denoising]. At training time, a sample from the dataset $x\sim\Omega$ is noised to a random step $t$ in the noising process, producing a partially noised sample $x_{t}\sim\mathcal{N}(x_{t-1}\sqrt{1-\beta_{t}},\beta_{t}I)$. The UNet is trained to produce $x_{t-1}$ from $x_{t}$ with the parameter $t$ given as a positional embedding. At inference time, a normal sample $x_{T}=z\sim\mathcal{N}(0,1)$ is produced and fed to the UNet, which reverses the noising steps one at a time and produces a new sample $x^{\prime}\sim\Omega^{\prime}$.
@@ 3.3.1 VAE Interpretation
Broadly speaking, there are two schools of thought by which to characterize the theory of diffusion models. We will begin with the VAE interpretation, which DDPM also subscribes to. Adapting the minimization problem (1) from the VAE section above gives
@@
Here $x_{t}$ represents the state of the datapoint $x_{0}$ at timestep $t$ in the Markov chain, which has $T$ total steps. $x_{T}$ is equivalent to the latent variable $z$, the fully noised datapoint, assumed to be drawn from a standard unit normal with dimension equal to that of the datapoint. The particular implementation used in this paper, denoising diffusion probabilistic models (DDPM) [ho2020denoising], parametrizes the forward noising process with a normal distribution
$q(x_{t}|x_{t-1})=\mathcal{N}(x_{t};\,x_{t-1}\sqrt{1-\beta_{t}},\beta_{t}I).$ (7)
The $\{\beta_{t}\}$ represent a fixed noise schedule, often linear in $t$, which controls the amount of noise added at each timestep and is chosen so that $q(x_{T}|x_{0})\approx\mathcal{N}(0,I)$ [song2021scorebased]. This gives rise to the backwards model, notably also a normal distribution, with analytical form
@@
We can do one better and give the forms of $\tilde{\mu}_{t}(x_{t},x_{0})$ and $\tilde{\beta}_{t}$. The full derivation can be found in [kong2021diffwaveversatile],
@@
and make the choice that $\Sigma_{\theta}(x_{t},t)=\sigma_{t}^{2}I$. We are finally ready to derive the ELBO, as we did with the VAE. Inserting into the minimization problem 6, and noting that the form of the term in inside the sum can be written [kong2021diffwaveversatile, sohldickstein2015deep, ho2020denoising]
@@
for some $C$ independent of the parameters $\theta$. By itself, this is suitable for optimization, but we can go one step further, using the reparameterization trick mentioned in 3.1 [kingma2013]. We separate the noise from the data to get
@@
for our minimization objective. In practice, we also drop the prefactor since it only scales the loss. In [ho2020denoising] it is noted that, aside from being simpler, this empirically leads to better results.
@@ 3.3.2 Continuous Limit
$x_{t}=\sqrt{1-\beta_{t}}x_{t-1}+\sqrt{\beta_{t}}\epsilon_{t},\,\,\,\,1\leq t\leq T$ (16)
for some $\epsilon_{t}$}, each independently identically distributed by the unit normal. It is natural to ask what happens in the continuous limit, namely as $T$ becomes very large. [song2021scorebased, 2023diff_survey, mcallester2023maths_diff] show that we arrive at the stochastic differential equation (SDE)
$dX_{t}=-\frac{1}{2}\beta(t)X_{t}dt+\sqrt{\beta(t)}dW_{t},$ (17)
where $\beta(t)$ is an appropriate continuation of the discrete noise schedule, and $W_{t}$ is the standard Wiener process, i.e., a continuous-time stochastic process with independent Gaussian increments, modeling the accumulation of noise over time. The reverse is also true, as discretizing this equation again returns us to the case above [mcallester2023maths_diff]. In this sense SDEs can be seen as an overarching framework for diffusion models. Given that the noise schedule increases sufficiently quickly so that $x_{T}\approx\mathcal{N}(0,I)$, the time reversed (see the main theorem in [anderson1982reverse_sde]) SDE is then
this is known as a ‘variance preserving’ (VP) SDE. The probability distribution $p(\cdot;x_{t})$ (here $t$ is a continuous parameter) satisfies the Fokker-Planck equation, also known as Kolmogorov’s forwards equation, which is an ODE describing how the probability density decays with time [song2021scorebased]. This equation, like equation 18, can be used to sample from the original data distribution [2023diff_survey].
3.3.3 Score-Matching
Another possible interpretation of diffusion models is to understand them by analogy with score-matching methods, which learn the ‘score’ (the gradient of log probability) with methods such as Langevin dynamics [song2020generativemodelingestimatinggradients]. Note that the score is known for a Gaussian. Indeed, in the formalism we have used so far,
$\nabla\ln{\mathcal{N}(x;\,\mu,\sigma^{2})}=\frac{\mu-x}{\sigma^{2}}$ (19)
Apply the reparameterization trick from 3.1 to the distribution of $x_{t}$ and obtain
Learning this score function with mean-squared error leads to the same optimization problem as 15. Philosophically, this is much the same as models formulated under the VAE interpretation. Where the schools of thought differ is in the noising process (which in turns leads to a different sampling process). SBM (score-based matching) models utilize
$x_{t}=x_{t-1}+\sqrt{\sigma^{2}_{t}-\sigma^{2}_{t-1}}\epsilon_{t}\,\,\,\,1\leq t\leq T$ (21)
for some noise scales $\{\sigma_{t}\}$ typically set as a geometric sequence, which again ensure in the limit that the fully noised distribution is equivalent to $\mathcal{N}(0,I).$ Performing the limit procedure above on this process also leads to a SDE, known as a ‘variance exploding’ (VE) SDE, a careful treatment of which can be found in [song2021scorebased].
4 Dataset
4.1 Setup of Simulations and Measurements
We investigate different test cases, comprising a high-resolution dataset obtained from Large Eddy Simulation (LES) representing the flow around a cylinder, and lower-resolution datasets from experimental Particle Image Velocimetry (PIV) capturing the flow behind arrays of seven cylinders, which are described in the following.
4.1.1 Flow around a cylinder
(a) Flow around a cylinder
(b) A schematic sketch of the experiments for flow behind cylinder arrays.
(c) Flow behind array of seven cylinders in two different configuration of 2V8H, and 4V2H from left to right, respectively. For clarity, only one out of every 256 vectors is displayed.
Figure 4: Numerical domain (a) and field of view for the investigated test cases numerical and experimental test case, respectively. Flow fields for the cylinder arrays (c).
First, we investigate the flow around a circular cylinder at a Reynolds number of 3900, a classical test case that exhibits complex fluid dynamic behavior due to the coexistence of laminar and turbulent regimes. This flow is characterized by vortex shedding, forming a Kármán vortex street in the cylinder’s wake [von1911mechanismus, Parnaudeau:2008, Norberg:1994, Ong:1996, Beaudan:1995, Kravchenko:2000]. The vortex street consists of a coherent system of vortices whose rotational axes are aligned with the cylinder axis. This case has been extensively studied both experimentally and numerically, making it suitable for validation and data-driven modeling.
The viscous fluid is governed by the incompressible Navier-Stokes equations (NSE), which originate from the conservation of mass and momentum. Under incompressible and isothermal conditions, the mass conservation reduces to the divergence-free condition of the velocity field, and the energy equation can be neglected. Let $\mathbf{u}=(u,v,w)$ denote the velocity vector. The NSE in Cartesian coordinates are then given by
$\nabla\cdot\mathbf{u}=0,\quad\frac{\partial\mathbf{u}}{\partial t}+(\mathbf{u}\cdot\nabla)\mathbf{u}=-\frac{1}{\rho}\nabla p+\nu\nabla^{2}\mathbf{u},$ (22)
where $\rho$ is the fluid density, $\nu$ the kinematic viscosity, and $p$ the pressure field. While ergodicity in turbulent flows is not guaranteed in general, it can reasonably be assumed for this setup [frisch1995turbulence].
Direct numerical simulation of the NSE for this test case is computationally infeasible due to the wide range of scales in turbulence. Consequently, Large-Eddy Simulation (LES) is employed using the commercial solver ANSYS Fluent [winhart2021large]. LES resolves the large, energy-containing eddies explicitly, while the effects of smaller subgrid-scale motions are modeled [sagaut2005large, Hirsch:2007]. This is achieved through spatial filtering, which decomposes a flow variable $\varphi_{t}(x)$ into a filtered (resolved) component $\overline{\varphi}_{t}(x)$ and a subgrid-scale component $\dot{\varphi}_{t}(x)$,
$\varphi_{t}(x)=\overline{\varphi}_{t}(x)+\dot{\varphi}_{t}(x),$ (23)
with the filter defined implicitly by the computational grid. The filtered NSE then take the form
$\nabla\cdot\overline{\mathbf{u}}=0,\quad\frac{\partial\overline{\mathbf{u}}}{\partial t}+(\overline{\mathbf{u}}\cdot\nabla)\overline{\mathbf{u}}=-\frac{1}{\overline{\rho}}\nabla\overline{p}+\frac{1}{\mathrm{Re}}\nabla^{2}\overline{\mathbf{u}},$ (24)
where $\mathrm{Re}=U_{\infty}D/\nu$ is the Reynolds number, $U_{\infty}$ the freestream velocity, and $D$ the cylinder diameter. The LES approach is particularly suitable because the large vortex structures carry most of the turbulent kinetic energy (TKE), while smaller structures can be assumed to be isotropic and homogeneous in the bulk flow according to Kolmogorov’s local isotropy assumption [Ferziger:2008, Kolmogorov:1991].
The computational setup consists of a grid with approximately 15 million cells and a time step chosen to maintain a CFL number on the order of unity. A schematic representation of the numerical domain is shown in Fig. 4(a). The simulation was performed for 25,000 time steps after the initial transients, corresponding to a physical time of roughly 1.45 seconds. Time integration is carried out using a non-iterative advancement scheme combined with a fractional-step method for pressure-velocity coupling. Advective fluxes are treated with a bounded central scheme to minimize numerical dissipation and prevent unphysical damping of small-scale turbulence [winhart2021large]. Snapshots of the flow field are sampled at a frequency of 68.9 Hz, with sufficient temporal separation to reduce correlation between frames.
For data-driven modeling, the transient LES velocity fields are post-processed to generate two-dimensional grayscale images representing the fluctuating velocity magnitude (see figure˜5). Let $V(\xi,t)=(V_{x},V_{y},V_{z})$ denote the local velocity vector. The grayscale intensity at a spatial location $\xi$ corresponds to the absolute deviation of the local velocity magnitude $c(\xi,t)=\sqrt{V_{x}^{2}+V_{y}^{2}+V_{z}^{2}}$ from its time average $\overline{c}(\xi)=\frac{1}{T}\int_{0}^{T}c(\xi,t)\,dt,$ i.e., $c^{\prime}(\xi,t)=|c(\xi,t)-\overline{c}(\xi)|.$
The fluctuation magnitude is then linearly scaled to an 8-bit intensity range $[0,255]$, such that darker regions correspond to nearly steady flow with weak fluctuations, whereas brighter regions mark locations of strong unsteady motion in the wake, providing a physically interpretable field of velocity fluctuation strength. This projection mapping preserves ergodicity in the reduced state space. The dataset contains $100,000$ images at $1,000\times 600$ resolution and is available at [winhart2024data]. The transient velocity fields were mapped onto structured evaluation grids, which define the sampling points for the images. Given the numerical domain length of $20D$ (see figure˜4(a)) and an image resolution of $1,000$ columns in the streamwise direction, one image column corresponds to a physical distance of $0.02D$. Because the evaluation grids were adapted to the required spatial resolution, the image sampling does not introduce scales smaller than those resolved by LES, ensuring that the image resolution is consistent with the LES filter scale and reflects the number of physically resolved degrees of freedom [winhart2021large].
4.1.2 Flow behind arrays of seven cylinders
The flow behind two arrays of seven cylinders each with different configurations was chosen, as although the Reynolds number based on a single cylinder is only $Re\sim 100$, it is already unsteady. The arrangement of multiple cylinders in the flow results in multiple individual wake flows that interact with each other. The vortex shedding in the individual wake flows does not necessarily show the same frequency, resulting in very broad and rich dynamics depending on the vertical and horizontal distances between the cylinders. If they are very close, they act as a single bluff body (where one may redefine the Reynolds number). On the other end, if they are far apart, the resulting flow can be characterized as co-shedding with specific interactions between the individual wakes that might differ in frequency and phase depending further on the distances.
The most complex case is the transitional state where the distances between the cylinders are large enough to prevent the array from acting as a single bluff body and still not large enough to let each cylinder develop its individual wake. As a result, irregular and asymmetric flow patterns arise behind the array, with varying vortex shedding frequencies over space and time. For further details on the flow physics, the interested reader is referred to [SharifiGhazijahani2023]. For this reason, the datasets 2V8H and 4V2H were chosen for further processing. The cylinders were arranged in two columns: four in the front and three in the rear, where the horizontal distance between these two columns (H) and the vertical distance between cylinders inside each column (V) are given in cylinder diameters.